A bullet of mass 10 g is fired horizontally into a block of wood of mass 2 kg and suspended like a ballistic pendulum. The bullet sticks in the block and the impact causes the block to swing so that its center of gravity rises 0.1 m. Find the velocity of the bullet just before the impact.

Can someone also explain the steps to me? I am very confused on all of this. Thank you
Answer #1

given,

mass of bullet = 10 gm

mass of wood = 2 kg

by conservation of energy

initial energy = final energy

0.5 * (2 + 10 * 10^-3) * v^2 = (2 + 10 * 10^-3) * g * 0.1

0.5 * v^2 = g * 0.1

v = 1.4 m/s

by conservation of momentum

initial momentum = final momentum

10 * 10^-3 * v = (2 + 10 * 10^-3) * 1.4

v = 281.4 m/s

velocity of bullet just before the impact = 281.4 m/s